If 9km distance can be cover in 4 hour,how much time needed to cover 24km
Answer by stanbon(75887)

If 9km distance can be cover in 4 hour,how much time needed to cover 24km
---------------
Use a proportion:
t/24 = 4/9
----
t = (24*4)/9 = 10 2/3 hrs
-------------------------------
Cheers,
Stan H.
---------------------